A container has 20 L of a mixture of alcohol and water which is 40% alcohol by volume. How much of the mixture should be removed and replaced by an equal volume of water so that the resulting solution will be 25% alcohol by volume?

Found 2 solutions by mananth, mathstutor494:
Answer by mananth(16946)

percent percentage ---------------- quantity
Alcohol 40 ---------------- x liters
water 0 ---------------- 20 - x liters
Mixture 25.00% ---------------- 20

40x+0(20-x) = 25 * 20

40 x = 500

/ 40
x=12.5 liters 40.00% Alcohol
7.5liters water
7.5 liters of mixture has to be replaced with water

Answer by mathstutor494(120)

Current composition of mixture......
Alcohol : 20*40/100 = 8 L
So water 20-8 : 12 L
New composition of mixture required....
Alcohol : 20*25/100 = 5 L
So Water : 15 L
So mixture to be removed and replaced by an equal volume of water = 3L